OpenStudy (anonymous):

Determine the quadrant when the terminal side of the angle lies according to the following conditions: sin (t) < 0, csc (t) > 0.

OpenStudy (anonymous):

How? \(csc (t)=\dfrac{1}{sin(t)}\) and 1>0 always, sin (t) <0, how can csc (t) >0 ??

OpenStudy (anonymous):

sorry. i cant help. i didn't write the problem and i dont understand either.

OpenStudy (michele_laino):

since: \[\csc t= \frac{ 1 }{ \sin t }\] so you are searching for a quadrant in which sin t>0 and sin t<0. Is it possible?